# Hilbert space methods for approximating multi-output latent variable Gaussian processes

### v2

@@ 4.5 Latent variable estimation
For all of the simulation scenarios discussed in Section 4.1, we evaluate the estimation capabilities of the different fitted models specified in Section 4.2. We show model performance in terms of absolute bias $|\text{Bias}(x_{post},x_{true})|$ of the latent posterior estimates $x_{post}$ with respect to their true values $x_{true}$ as well as $\text{SD}(x_{post})$ as a measure of posterior sharpness. Since mean-field VI is known to underestimate the posterior SD (Kucukelbir et al., , 2017), we avoid using a metric like RMSE that favors (too) narrow posteriors, thus providing misleading interpretations if used as a sole metric.
@@
We summarize our findings using the multi-level model (Bürkner, , 2017) described in supplementary materials Section A for both posterior bias and SD. The SE simulation scenario presented in Figure 5 shows that, in most cases, HSGPs have lower bias compared to exact GPs and VIGPs. The VIGPs consistently show the highest bias. Overall, for the SE data scenario, we see on average about 11% (for higher sample sizes) to about 18% (for lower sample sizes) decrease in posterior bias between VIGPs and HSGPs (with $M=22$ basis functions). Results for the posterior SD are more mixed, with VIGPs showing higher SDs for small sample sizes and about the same SDs for larger sample sizes.
@@
In Figure 6 we show the more challenging simulation scenario. Here, the data generating process had a higher amount of variability in length-scales across the output dimensions. We observe a stronger difference in posterior bias between VIGPs and HSGPs (in favor of HSGPs). We hypothesize that HSGPs are more strongly favored here as they allow for varying hyperparameters across dimensions, while the VIGPs (at least their Pyro implementation) only support constant hyperparameters.
@@
Among the HSGPs, higher number of basis functions ($M=26\text{ }\text{and}\text{ }30$ compared to $M=22$) result in lower posterior bias for both the simulation scenarios shown in Figures 5 and 6. Combining this result with their corresponding model calibration tests, we recommend using $M=30$ to achieve the highest accuracy for the latent variable estimates. For faster but reliable results, HSGPs with $M=22$ can be used in simpler cases like the SE data scenario. Furthermore, in the remaining scenarios (see Figures S10-S13 in supplementary materials Section E), HSGPs perform consistently better than their competitors. In all the scenarios, each of the fitted models (exact GPs only for $N=20$) show a consistent decrease in posterior bias and SD as the number of output dimensions $D$ and sample size $N$ increases. We report the estimation results for GP hyperparameters in the supplementary materials Section B.
@@ 5 Real-World Case Study
We showcase the application of the HSGPs to real-world scRNA sequencing data by re-analyzing cell-cycle data from Mahdessian et al., (2021). Our aim is to estimate the underlying cell ordering known as ”pseudotime”. The dataset comprises of spliced single-cell RNA expression profiles along the cell cycle with their own time sequence for cells that are aligned with the experimental time called ”cell hours”. We choose this dataset because it covers a cyclic process going through three distinct biological phases (G1, G2M and S-ph) depicting substantial variation in gene expressions. Additionally, the provided cell hours is a continuous temporal sequence that can be utilized as a prior guide to estimate pseudotime. For the purpose of this case study, we use the data from all the cells along with a selection of 12 influential genes (see supplementary materials Section G for the list) based on the recommendations of Mahdessian et al., (2021). In other words, each sample point corresponds to a single cell and each output dimension corresponds to a single gene, with the value being the gene expressions per cell. Thus, the analyzed data comprises of $N=960$ observations and $D=12$ output dimensions.
Figure 7: (a) Difference of HSGP pseudotime estimates from the prior cell hours; (b) Difference of VIGP pseudotime estimates from the prior cell hours; (c) Difference of HSGP and VIGP pseudotime estimates. Both models assume an SE covariance function.
@@
We fit a HSGP and VIGP with SE covariance function. We specify the priors for length-scale $\rho\sim\text{Normal}^{+}(0.4,0.1^{2})$, GP marginal SD $\alpha\sim\text{Normal}^{+}(14,3.5^{2})$, and error SD $\sigma\sim\text{Normal}^{+}(7,3.5^{2})$. The prior specifications are based on the similar case study in Mukherjee et al., (2025). Based on the $\rho$ prior and the $(0,1)$ range of the input space, we use the minimum number of basis function $M=6$ (see Equation 25). For the VIGP, we again resort to default priors since custom priors are not supported. Similarly, we use 10 inducing points as the reasonable choice without compromising the numerical stability.
@@
In Figure 8, we show the posterior estimates of the HSGP hyperparameters along with their posterior 95% CIs. A key observation here is the strong differences in the hyperparameter estimates across the genes (output dimensions). This confirms how crucial it is for the model to assume gene-specific hyperparameters while modeling these biological processes. Modeling varying GP hyperparameters, while supported in our HSGPs, is not possible yet for the (pyro implemented) VIGPs, so we do not include their estimates in Figure 8.
The runtime for the HSGP was approximately 16 minutes on an Apple M4 chip (2 cores in parallel) with 24 GB working memory. Fitting the VIGP took about 3 minutes using the cluster computing resources detailed in Section 4.2.
@@ 6 Discussion
We generalize Hilbert space approximations for Gaussian processes (HSGPs; Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022) to develop a scalable class of latent variable models for high-dimensional multi-output data. Our HSGPs account for both varying hyperparameters across output dimensions and correlations between output dimensions. They reduce the steep computational complexity of $O(N^{3}D+ND^{2})$ for exact multi-output GPs to $O(NMD+ND^{2})$ for a dataset of $N$ samples and $D$ output dimensions, where $M$ is the number of basis functions of the HSGP. Since $M$ is typically much smaller than $N$, using HSGPs reduce computation times drastically. Through extensive simulations, we investigated the statistical properties of latent GPs in terms of model convergence, uncertainty calibration, and estimation accuracy. In all of the above criteria, the HSGPs performed better than both exact GPs and approximate GPs based on inducing points and variational inference (VIGPs)(Titsias, , 2009; Ahmed et al., , 2019). While VIGPs are faster yet, they provide limited modeling capabilities and are outperformed by HSGPs in terms of model calibration and estimation accuracy. Even though HSGPs are comparably slower, they only need a few minutes even for relatively large datasets such as the one investigated in our real-world case study. This stands in sharp contrast to exact GPs, which would have taken at least several hours, if not more, to fit the same data.
@@ 6.1 Limitations and future research
With regard to the correlation structure $C$ across output dimensions (see Section 3.3.1), we make the choice of modeling it on the level of the GP functions $f$ rather than on the level of observations $y$. It is not always apparent which of these two options is preferable for a given application. For example, in spatial statistics, correlations between discrete locations are also either modeled on the mean (similar to the GP functions in our case) level, e.g., via conditional autoregressive (CAR) structures, or on the observation level, e.g., via spatial autoregressive (SAR) structures (Wall, , 2004). In this paper, we apply our methods to RNA gene expressions for cell cycle data, where the genes are known to be correlated at the biological process level (Riba et al., , 2022). Thus, specifying across-dimension correlations on the GP functions appear sensible here. For other applications, it may be preferable to specify the correlations at the observational level, or even jointly at both levels.
Another consideration concerning $C$ is whether to model it as input dependent or input independent. In this paper, we chose it to be input independent, thus estimating the full correlation matrix directly. As an alternative, an input dependent formulation could be achieved through another GP that models the correlations across output dimensions within each observation (similar to a cross-covariance structure shown in Gneiting et al., (2010)). In this regard, it would further be interesting to study Hilbert space approximations also for this across-output dimension GP, especially when the number of dimensions is large enough for exact GPs to become computationally impractical. We leave these extensions of our framework to future research.
Lastly, an important generalization of latent variable GPs concern the inclusion of derivative information, which can further increase estimation accuracy (Mukherjee et al., , 2025). However, such methods have so far only been developed for exact GPs, which hamper their real-world applicability. Extending HSGPs to include derivative information will likely provide a scalable solution. However, such an extension requires spectral densities for composite covariance functions, which to our knowledge is yet to be derived.
@@ Acknowledgments
We acknowledge the valuable insights provided by members of ClaassenLab and BürknerLab. Paul-Christian Bürkner further acknowledges support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Soham Mukherjee.
@@
To analyse the results from our experiments, we use a multilevel analysis of variance model (ANOVA) fitted with brms (Bürkner, , 2017), which disentangles the various components of our simulation study design. With $\mu_{resp}$ and $\sigma_{resp}$ being the mean and SD of our response variable, we use
@@
where $\beta,\text{ }b$ (for $\mu_{resp}$) and $\eta,\text{ }u$ (for $\sigma_{resp}$) are coefficients at the population and group levels with $X$ and $Z$ being the corresponding design matrices. The $s_{\tau}(t_{\tau})$ terms denote smooth functions over covariates $t$ fitted via splines. We use these models to summarize the results from our experiments. In the simulation studies, the population level design matrix $X$ contains covariates representing the approximation method and number of output dimensions along with their interaction terms. Based on our experimental setup, the approximation methods include the exact model, HSGPs (with three choices of basis functions) and VIGPs, thus creating a five level factor variable. We use a three level factor variable depicting the $5$, 1010, and $20$ output dimensions. Through $Z$ we account for the group-level dependency structure in the response induced by fitting multiple models to the same simulated dataset. We include a random intercept over datasets as well as corresponding random slopes of the approximation method choices. Lastly, to capture the non-linear relation of the response to the ground truth $t$, we introduce thin-plate regression spline terms $s_{\mu_{\tau}}(t_{\tau})$ and $s_{\sigma_{\tau}}(t_{\tau})$. The spline terms accounts for any non-linear relationship of the response with respect to the true parameter values $t$. We use the log $\gamma$ scores as our response to summarize the model calibration results in Section 4.4. To summarize the results of latent variable estimation in Section 4.5, we subsequently specify posterior bias and SD as the response.
@@
We check SBC by computing the confidence bands for the empirical cumulative distribution function (ECDF) of the rank distribution and visualising it under the assumption of the uniformity. An alternative, however is based on the probability of $\gamma$ observing the most extreme point on the ECDF under the assumption of uniformity. The test statistic is given by
@@
We present the RMSE for each model hyperparameter under the SE data scenario in Figure 9 where we see how the decreasing effect on the RMSE of GP marginal SD $\alpha$ and error SD $\sigma$ as the sample size $N$ increases. Since the results were qualitatively same for all the different choices of output dimensions $D$, we decided to showcase only for the $D=20$. Interestingly, RMSE for the length-scale $\rho$ stays similar across all sample size choices. We can thus be sure about the reliability of HSGPs in recovering true hyperparameter values for all kinds of data scenarios. In the more challenging case of the highly variable length-scale, we see in Figure 10 that the RMSE for the length-scale $\rho$ increases as expected, but is still comparatively better than for the exact GP for $N=20$. All the other simulation study scenarios are presented as additional plots in the Appendix F (see Figures 22-25).
@@
Figure 11: Matern 3/2 scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs and HSGPs. The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 12: Matern 5/2 scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs and HSGPs. The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 13: Periodic data with lower oscillation scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs and HSGPs. The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 14: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs and HSGPs. The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 15: Squared exponential scenario (highly varying $\rho$): Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs and HSGPs. The y-axes for Bulk and Tail ESS plots are log10 transformed.
D: Model calibration tests (additional figures)
@@
Figure 16: Matern 3/2 scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behaviour of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
Figure 17: Matern 5/2 scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behaviour of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
E: Latent input estimation (additional figures)

### v3

@@ 4.5 Latent variable estimation
For all of the simulation scenarios discussed in Section 4.1, we evaluate the estimation capabilities of the different fitted models specified in Section 4.2. We show model performance in terms of absolute bias $|\text{Bias}(x_{post},x_{true})|$ of the latent posterior estimates $x_{post}$ with respect to their true values $x_{true}$ as well as $\text{SD}(x_{post})$ as a measure of posterior sharpness. Since mean-field VI is known to underestimate the posterior SD (Kucukelbir et al., 2017), we avoid using a metric like RMSE that favors (too) narrow posteriors, thus providing misleading interpretations if used as a sole metric.
@@
We summarize our findings using the multi-level model (Bürkner, 2017) described in supplementary materials Section A for both posterior bias and SD. The SE simulation scenario presented in Figure 5 shows that, in most cases, HSGPs have lower bias compared to exact GPs and VIGPs. The VIGPs consistently show the highest bias. Overall, for the SE data scenario, we see on average about 11% (for higher sample sizes) to about 18% (for lower sample sizes) decrease in posterior bias between VIGPs and HSGPs (with $M=22$ basis functions). Results for the posterior SD are more mixed, with VIGPs showing higher SDs for small sample sizes and about the same SDs for larger sample sizes. Specifically, HSGPs overcome the computational challenges (in MCMC convergence and posterior uncertainty calibration) of exact GPs to provide better latent variable estimation accuracy.
@@
In Figure 6 we show the more challenging simulation scenario. Here, the data generating process had a higher amount of variability in length-scales across the output dimensions. We observe a stronger difference in posterior bias between VIGPs and HSGPs (in favor of HSGPs). We hypothesize that HSGPs are more strongly favored here as they allow for varying hyperparameters across dimensions, while the VIGPs (at least their Pyro implementation) only support single hyperparameters across all output dimensions. For example, a single length scale parameter $\rho$ is optimized for all $D$ output dimensions. While this is a limitation of the current VIGP implementation, we compare a favorable case where the GP hyperparameters are fixed to the data generating conditions in the special simulation scenario with $N=1000$ in Section 4.6.
@@
Among the HSGPs, higher number of basis functions ($M=26\text{ }\text{and}\text{ }30$ compared to $M=22$) result in lower posterior bias for both the simulation scenarios shown in Figures 5 and 6. Combining this result with their corresponding model calibration tests, we recommend using $M=30$ to achieve the highest accuracy for the latent variable estimates. For faster but reliable results, HSGPs with $M=22$ can be used in simpler cases like the SE data scenario. Furthermore, in the remaining scenarios (see Figures S16-S19 in supplementary materials Section F), HSGPs perform consistently better than their competitors. In all the scenarios, each of the fitted models (exact GPs only for $N=20$) show a consistent decrease in posterior bias and SD as the number of output dimensions $D$ and sample size $N$ increases. We report the estimation results for GP hyperparameters in the supplementary materials Section B (followed by additional figures in Section H).
4.6 Checking robustness of HSGPs with wide priors
In the above discussed GP simulation scenarios, the hyperparameter priors for exact GPs and HSGPs are aligned with the data generating conditions as a requirement for SBC checks (see Section 4.4). Therefore, in those simulation studies, the exact GPs and HSGPs receive a favorable treatment as compared to VIGPs. Hence, to check the robustness of latent variable HSGPs, we perform additional simulations where we apply hyperparameter priors to the HSGPs that are less informed (i.e., wider priors) along with a larger sample size $N=1000$. We leave exact GPs out of this simulation study due to their computational limitations.
Keeping the data generating conditions same as discussed in Section 4.1, we specify priors for length-scale $\rho\sim\text{Normal}^{+}(1.5,0.5^{2})$, GP marginal SD $\alpha\sim\text{Normal}^{+}(3.5,1^{2})$ and error SD $\sigma\sim\text{Normal}^{+}(1.5,1^{2})$. That is, we shift the prior mean away from the data generating conditions along with a much wider prior variance. A comparison of the data generating conditions and the GP hyperparameter priors are presented in Table 2. Simulations are carried out for SE, Matern 3/2 and 5/2 covariance functions. Based on the specified length-scale prior and chosen covariance function, the minimum required number of basis functions for HSGPs change according to the empirical relation shown in Eq.(25)-(27).
Table 2: Comparison of GP hyperparameter priors and data generating conditions
Hyperparameter Dgp prior Model prior
$\rho$ $\text{Normal}^{+}(1,0.05^{2})$ $\text{Normal}^{+}(1.5,0.5^{2})$
$\alpha$ $\text{Normal}^{+}(3,0.25^{2})$ $\text{Normal}^{+}(3.5,1^{2})$
$\sigma$ $\text{Normal}^{+}(1,0.25^{2})$ $\text{Normal}^{+}(1.5,1^{2})$
  • Note: The Dgp priors show the true sampling distribution for simulated data and Model priors show the prior distribution for the HSGP parameters.
We consider two VIGP variants: VIGP (default) and VIGP (prior). The former is the same VIGP model considered in other simulation scenarios with the default priors used by Pyro. In contrast, for VIGP (prior), we fix the GP hyperparameters to the mean of the true sampling distributions, thereby acting as a highly informed reference.
Figure 7: Squared exponential scenario (special case): posterior bias and SD on recovery of latent inputs for all fitted models. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions. VIGP (default) shows the default Pyro implementation and VIGP (prior) has fixed GP hyperparameters informed by the true sampling distributions.
We summarize the posterior bias and SD using the same multilevel model (described in supplementary materials Section A) used for other simulation scenarios. For the SE scenario in Figure 7, the HSGPs achieve lower bias compared to VIGPs with a higher number of basis functions $M$. This is especially true for higher output dimensions $D=20$ where we need much higher $M$ than the minimum $M$ suggested by the empirical relation in Eq.(25). On the other hand, posterior SDs for HSGPs are consistently lower than VIGPs. Among the VIGP variants, both VIGP (default) and VIGP (prior) performs similarly in terms of latent variable estimation accuracy. That is, fixing the GP hyperparameters for VIGPs to informed values does not result in any improvement in model performance. The results for Matern 3/2 and 5/2 covariance functions (see Figures S20 and S21 in supplementary materials section G) are highly similar, with Matern 3/2 showing even more pronounced differences between HSGPs and VIGPs.
The average runtime per dataset of sample size $N=1000$ on the computing cluster (details mentioned previously) was about 6.6 hours on average for SE, 11.5 hours for Matern 3/2 and 10.2 hours for Matern 5/2 models. For these specific simulation conditions, the runtime for HSGPs are reasonable compared to exact GPs that are computationally infeasible. Although the HSGP computation times are affected due to wider prior specifications for GP hyperparameters, the model performance in terms of latent variable estimation accuracy remains superior to VIGPs as in other simulation scenarios. In terms of speed, the VIGPs were similarly fast as in other simulation scenarios.
@@ 5 Real-World Case Study
We illustrate the application of the HSGPs using a real-world scRNA sequencing data. To that end, we re-analyze cell-cycle data from Mahdessian et al. (2021). Our aim is to estimate the underlying cell ordering known as ”pseudotime”. The dataset comprises of spliced single-cell messenger RNA (or mRNA) expression profiles of cells along the cell cycle. Each cell is additionally labeled with its own approximated time in the cell cycle process, i.e. the experimental time called ”cell hours”. We choose this dataset because it covers a cyclic process going through three distinct biological phases (G1, G2M and S-ph) depicting substantial variation in gene expressions. Additionally, the provided cell hours is a continuous temporal sequence that can be utilized as a prior guide to estimate pseudotime. For the purpose of this real-world case study, we use the data from all the cells along with a selection of 12 influential genes (see supplementary materials Section I for the list) based on the recommendations of Mahdessian et al. (2021). In other words, each sample point corresponds to a single cell and each output dimension corresponds to a single gene, with the value being the gene expressions per cell. Thus, the analyzed data comprises of $N=960$ observations and $D=12$ output dimensions.
Figure 8: (a) Difference of HSGP pseudotime estimates from the prior cell hours; (b) Difference of VIGP pseudotime estimates from the prior cell hours; (c) Difference of HSGP and VIGP pseudotime estimates. Both models assume an SE covariance function. The different colors denote the three distinct biological phases (G1, G2M and S-ph).
@@
We fit a HSGP and VIGP with SE covariance function. We specify the priors for length-scale $\rho\sim\text{Normal}^{+}(0.4,0.1^{2})$, GP marginal SD $\alpha\sim\text{Normal}^{+}(14,3.5^{2})$, and error SD $\sigma\sim\text{Normal}^{+}(7,3.5^{2})$. The prior specifications are based on the similar real-world case study in Mukherjee et al. (2025). Based on the $\rho$ prior and the $(0,1)$ range of the input space, we use the minimum number of basis function $M=6$ (see Eq.(25)). For the VIGP, we again resort to default priors since custom priors are not supported. Additionally, based on the $N=1000$ simulation study results in Section 4.6, the default implementation performs equally well compared to fixed hyperparameter settings. For the number of inducing points, we follow Ahmed et al. (2019) where $10$ inducing points are shown to be sufficient for pseudotime estimation. Moreover, keeping the number of inducing points to $10$ similar to our simulation studies allows us to avoid compromising the numerical stability of VIGPs.
@@
In Figure 9, we show the posterior estimates of the HSGP hyperparameters along with their posterior 95% CIs. A key observation here is the strong differences in the hyperparameter estimates across the genes (output dimensions). This confirms the importance of assuming gene-specific hyperparameters while modeling these biological processes. Modeling varying GP hyperparameters, while supported in our HSGPs, is not possible yet for the (pyro implemented) VIGPs, so we do not include their estimates in Figure 9.
The runtime for the HSGP was approximately 16 minutes on an Apple M4 chip (2 cores in parallel) with 24 GB working memory. Same model fitted using the cluster computing resources detailed in Section 4.2 took about 58 minutes. In comparison, fitting the VIGP took about 3 minutes using the same cluster computing resources.
@@ 6 Discussion
We generalize Hilbert space approximations for Gaussian processes (HSGPs; Solin and Särkkä, 2020; Riutort-Mayol et al., 2022) to develop a scalable class of latent variable models for high-dimensional multi-output data. Our HSGPs account for both varying hyperparameters across output dimensions and correlations between output dimensions. They reduce the steep computational complexity of $O(N^{3}D+ND^{2})$ for exact multi-output GPs to $O(NMD+ND^{2})$ for a dataset of $N$ samples and $D$ output dimensions, where $M$ is the number of basis functions of the HSGP. Since $M$ is typically much smaller than $N$, using HSGPs reduce computation times drastically. Through extensive simulations, we investigated the statistical properties of latent GPs in terms of model convergence, uncertainty calibration, and estimation accuracy. In all of the above criteria, the HSGPs performed better than both exact GPs and approximate GPs based on inducing points and variational inference (VIGPs)(Titsias, 2009; Ahmed et al., 2019). While VIGPs are faster yet, they provide limited modeling capabilities and are outperformed by HSGPs in terms of model calibration and estimation accuracy. Moreover, even with less informed hyperparameter priors (wide priors), the HSGPs achieve higher accuracy than VIGPs. Although HSGPs are comparably slower, they are reasonably fast even for relatively large datasets such as the one investigated in our real-world case study. This stands in sharp contrast to exact GPs, which would have taken at least several hours, if not more, to fit the same data.
@@ 6.1 Limitations and future research
With regard to the correlation structure $C$ across output dimensions (see Section 3.3.1), we make the choice of modeling it on the level of the GP functions $f$ rather than on the level of observations $y$. It is not always apparent which of these two options is preferable for a given application. For example, in spatial statistics, correlations between discrete locations are also either modeled on the mean (similar to the GP functions in our case) level, e.g., via conditional autoregressive (CAR) structures, or on the observation level, e.g., via spatial autoregressive (SAR) structures (Wall, 2004). In this paper, we apply our methods to RNA gene expressions for cell cycle data, where the genes are known to be correlated at the biological process level (Riba et al., 2022). Thus, specifying across-dimension correlations on the GP functions appear sensible here. For other applications, it may be preferable to specify the correlations at the observational level, or even jointly at both levels.
Another consideration concerning $C$ is whether to model it as input dependent or input independent. In this paper, we chose it to be input independent, thus estimating the full correlation matrix directly. As an alternative, an input dependent formulation could be achieved through another GP that models the correlations across output dimensions within each observation (similar to a cross-covariance structure shown in Gneiting et al. (2010)). In this regard, it would further be interesting to study Hilbert space approximations also for this across-output dimension GP, especially when the number of dimensions is large enough for exact GPs to become computationally impractical. We leave these extensions of our framework to future research.
Lastly, an important generalization of latent variable GPs concern the inclusion of derivative information, which can further increase estimation accuracy (Mukherjee et al., 2025). However, such methods have so far only been developed for exact GPs, which hamper their real-world applicability. Extending HSGPs to include derivative information will likely provide a scalable solution. However, such an extension requires spectral densities for composite covariance functions, which to our knowledge is yet to be derived.
@@ Acknowledgments
We acknowledge the valuable insights provided by all group members. P.B further acknowledges support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting S.M.
@@
To analyze the results from our experiments, we use a multilevel analysis of variance model (ANOVA) fitted with brms (Bürkner, 2017), which disentangles the various components of our simulation study design. With $\mu_{resp}$ and $\sigma_{resp}$ being the mean and SD of our response variable, we use
@@
where $\beta,\text{ }b$ (for $\mu_{resp}$) and $\eta,\text{ }u$ (for $\sigma_{resp}$) are coefficients at the population and group levels with $X$ and $Z$ being the corresponding design matrices. The $s_{\tau}(t_{\tau})$ terms denote smooth functions over covariates $t$ fitted via splines. We use these models to summarize the results from our experiments. In the simulation studies, the population level design matrix $X$ contains covariates representing the approximation method and number of output dimensions along with their interaction terms. Based on our experimental setup, the approximation methods include the exact model, HSGPs (with three choices of basis functions) and VIGPs, thus creating a five level factor variable. We use a three level factor variable depicting the $5$, 1010, and $20$ output dimensions. Through $Z$ we account for the group-level dependency structure in the response induced by fitting multiple models to the same simulated dataset. We include a random intercept over datasets as well as corresponding random slopes of the approximation method choices. Lastly, to capture the non-linear relation of the response to the ground truth $t$, we introduce thin-plate regression spline terms $s_{\mu_{\tau}}(t_{\tau})$ and $s_{\sigma_{\tau}}(t_{\tau})$. The spline terms accounts for any non-linear relationship of the response with respect to the true parameter values $t$. We use the log $\gamma$ scores as our response to summarize the model calibration results. To summarize the results of latent variable estimation, we subsequently specify posterior bias and SD as the response.
@@
We check SBC by computing the confidence bands for the empirical cumulative distribution function (ECDF) of the rank distribution and visualizing it under the assumption of the uniformity. An alternative, however is based on the probability of $\gamma$ observing the most extreme point on the ECDF under the assumption of uniformity. The test statistic is given by
@@
We present the RMSE for each model hyperparameter under the SE data scenario in Figure 10 where we see how the decreasing effect on the RMSE of GP marginal SD $\alpha$ and error SD $\sigma$ as the sample size $N$ increases. Since the results were qualitatively same for all the different choices of output dimensions $D$, we decided to showcase only for the $D=20$. Interestingly, RMSE for the length-scale $\rho$ stays similar across all sample size choices. We can thus be sure about the reliability of HSGPs in recovering true hyperparameter values for all kinds of data scenarios. In the more challenging case of the highly variable length-scale, we see in Figure 11 that the RMSE for the length-scale $\rho$ increases as expected, but is still comparatively better than for the exact GP for $N=20$. All the other simulation study scenarios are presented as additional plots in the Section F (see Figures 31-34).
@@
We present the model convergence diagnostics of exact GPs and HSGPs for all the simulation scenarios except for the squared exponential (SE) scenario (which is shown in the main manuscript). In all of the scenarios, the HSGPs consistently the 1.01 threshold of model convergence as suggested by Vehtari et al. (2021). HSGPs subsequently also have much higher Bulk and Tail-ESS. Based on the diagnostics, HSGPs show much more consistent and stable convergence as compared to exact GPs.
@@
Figure 12: Matern 3/2 scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs (for $N=20$) and HSGPs (combined for $N=20,50\text{ }\text{and}\text{ }200$ cases). The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 13: Matern 5/2 scenario: Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs (for $N=20$) and HSGPs (combined for $N=20,50\text{ }\text{and}\text{ }200$ cases). The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 14: Periodic data scenario (lower oscillation): Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs (for $N=20$) and HSGPs (combined for $N=20,50\text{ }\text{and}\text{ }200$ cases). The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 15: Periodic data scenario (higher oscillation): Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs (for $N=20$) and HSGPs (combined for $N=20,50\text{ }\text{and}\text{ }200$ cases). The y-axes for Bulk and Tail ESS plots are log10 transformed.
@@
Figure 16: Squared exponential scenario (highly varying $\rho$): Convergence check for (a) latent inputs and (b) GP hyperparameters of the exact GPs (for $N=20$) and HSGPs (combined for $N=20,50\text{ }\text{and}\text{ }200$ cases). The y-axes for Bulk and Tail ESS plots are log10 transformed.
D: Further MCMC convergence analysis
We further show MCMC convergence comparison for latent $x$, length-scale $\rho$, GP marginal SD $\alpha$ and error SD $\sigma$ between exact GPs and HSGPs for $N=20$ case. Overall, HSGPs promise superior MCMC convergence compared to exact GPs, especially for latent variable inputs. Further, exact GPs struggle with convergence especially for smaller latent inputs as well as smaller length-scales as shown in the right panel of Figures 17 and 18 respectively. As a consequence of these computational issues, exact GPs show miscalibrations in SBC checks and thus lower latent variable estimation accuracy compared to HSGPs.
Figure 17: Squared exponential scenario: MCMC convergence comparison between exact GPs and HSGPs ($N=20$ case) for latent inputs through (a) Rhat, (b) Bulk-ESS and (c) Tail-ESS measures. The behavior of each measures across ground truth are shown in the right-hand panel.
Figure 18: Squared exponential scenario: MCMC convergence comparison between exact GPs and HSGPs ($N=20$ case) for length-scale through (a) Rhat, (b) Bulk-ESS and (c) Tail-ESS measures. The behavior of each measures across true values are shown in the right-hand panel.
Figure 19: Squared exponential scenario: MCMC convergence comparison between exact GPs and HSGPs ($N=20$ case) for GP marginal SD through (a) Rhat, (b) Bulk-ESS and (c) Tail-ESS measures. The behavior of each measures across ground truth are shown in the right-hand panel.
Figure 20: Squared exponential scenario: MCMC convergence comparison between exact GPs and HSGPs ($N=20$ case) for error SD through (a) Rhat, (b) Bulk-ESS and (c) Tail-ESS measures. The behavior of each measures across ground truth are shown in the right-hand panel.
E: Model calibration tests (additional figures)
We present the ECDF plots for the fitted models in the Matern 3/2 and 5/2 simulation scenario with $N=20$ and $D=20$. Similar issues with calibration as in SE case presented in the main paper is seen with exact GPs and VIGPs.
We present the log $\gamma$ scores (offset by the 95% confidence threshold) for all the fitted models in the Matern 3/2 and 5/2 simulation scenarios. The exact GPs and VIGPs consistently fail the calibration test for $D=20$. The HSGPs with $M=51$ number of representative basis functions show good calibration across all the scenarios.
@@
Figure 23: Matern 3/2 scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity, that is, models with values less than 0 are miscalibrated. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
Figure 24: Matern 5/2 scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity, that is, models with values less than 0 are miscalibrated. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
F: Latent input estimation (additional figures)
We present the model evaluation on latent variable estimation accuracy for the Matern 3/2, 5/2 as well as the periodic data (both high and low oscillation) scenarios. In almost all of the cases, HSGPs have lower bias compared to exact GPs and VIGPs. In case of posterior SDs, the results are mixed. VIGPs sometimes show lower posterior SD as compared to HSGPs whereas in other cases, they are higher or same. Combined with the calibration test results, VIGPs are likely to underestimate posterior uncertainty due to severe model miscalibrations.